# How Post-Training Reshapes LLMs: A Mechanistic View on Knowledge, Truthfulness, Refusal, and Confidence

### v2

@@ 4 Knowledge Storage and Representation
LLMs are known to store factual knowledge in their parameters, particularly in “knowledge neurons” and MLP layers that act as key-value memories. This enables them to answer factual queries, such as answering “TRUE” or “FALSE” for prompt “The city of New York is in the United States. This statement is:”. While such knowledge is believed to emerge during pre-training and persist through post-training, mechanistic evidence remains limited. As knowledge is foundational for LLMs, we first examine how post-training affects it—whether it alters (1) knowledge-storage locations and (2) knowledge representations.
When prompted to classify a statement’s truthfulness, LLMs retrieve stored knowledge into hidden representations at some layers and tokens, which guide the final output. Following Marks & Tegmark (2024), we adapt causal tracing to identify knowledge-storage locations by patching hidden states between true and false statement pairs. Each pair is token-aligned and differs only in subject—e.g., “The city of Seattle is in France.” vs. “The city of Paris is in France.”. The relation (e.g., city-in-country) is true for only one statement. We treat subject and object tokens (e.g., city and country) as knowledge-related target tokens for analysis.
@@
where a high value indicates that some knowledge is stored in the $l$-th layer at the $i$-th token.
In order to aggregate the location of individual knowledge and analyze the knowledge storage location in general, we average the patching results over all the statements, where we carefully curate the statements to have the same token lengths and token positions. We use (true, false) statement pairs for patching, where each pair only differs in their subjects, and we explain the dataset construction details in Appendix B.1. We construct input prompts by 4-shot examples containing 2 true statements and 2 false statements, followed by the final statement. Patching is applied to the final statement using the methods described above. The aggregated results ($\tilde{M}_{i}^{l}$) are normalized ($M_{i}^{l}$) for better visualization:
$\tilde{M}_{i}^{l}=\frac{1}{|D|}\sum_{(s,\hat{s})\in D}M_{i}^{l}(s,\hat{s}),\quad M_{i}^{l}=normalize(\tilde{M}_{i}^{l})$ (2)
In normalization, we divide the range $\min_{i,l}\tilde{M}_{i}^{l},\max_{i,l}\tilde{M}_{i}^{l}]$ into 20 equal-width bins. We set values in the lower 10 bins to 0 and values in the upper 10 bins to 0.1, 0.2, …, 1. We denote the normalized result as $M_{model}\in R^{L*I}$, where $L$ and $I$ are the number of layers and tokens.
@@
Q1: Does post-training change LLM’s knowledge storage locations? Figure 2 visualizes the results ($M_{model}$) of Llama-3.1-8B base and instruct on the cities dataset. As shown in the left figure, influential patching consistently occurs at three token positions: subject, object, and the last token. Subject and object are important for both base and instruct. Their difference is nearly zero (e.g., (c)), indicating that base and instruct store knowledge in nearly identical locations. This pattern holds across all datasets and models, with additional visualizations in Appendix B.5. We further conduct quantitative analysis and include sft models. We compute Pearson correlation coefficient between $M_{\textsc{base}}$ and $M_{\textsc{post}}$, where post is instruct or sft. We also measure the maximum absolute difference value over all tokens, $max|M_{\textsc{post}}-M_{\textsc{base}}|$, as well as only over knowledge-related tokens (subject and object), $max|M_{\textsc{post}}-M_{\textsc{base}}|_{K}$. Results for the Llama-3.1-8B family are in Table 1, and for Mistral-7B in Table 8 in Appendix B.4. All results show high correlation and low difference, confirming that post-training has little influence on knowledge-storage locations.
@@
Q2: Does post-training change the knowledge representations? We further conduct cross-model experiments by patching hidden representations from base to post (forward patching) and from post to base (backward patching). It allows us to analyze whether knowledge representations in base can still work in post, and vice versa.
Due to space limits, we put the visualizations on all models and datasets in Appendix B.5. The results demonstrate that the forward patching is almost always successful, but the backward patching often fails, i.e., it does not recover the log probability difference. It leads to the conclusion that knowledge representations of base still work after post-training, but post-training also develops new knowledge representations.
Verification on in-distribution dataset One natural question is that our previous experiments are based on general datasets independent of post-training, which can be considered out-of-distribution. To verify the conclusions completely, we extract factual knowledge from the Tulu dataset (Lambert et al., 2024), which was used to fine-tune Llama-3.1-8B-SFT and Mistral-7B-v0.3-SFT (Feuer et al., 2025). We generate (true, false) statement pairs from the dataset, and it can be considered an in-distribution dataset for the sft models. Different from previous datasets, pairs in the Tulu dataset could have different lengths, so we slightly modify the metric calculation, specified in Appendix B.3. The last column of Table 1 shows results of the Llama-3.1-8B family, and the last column of Table 8 in Appendix B.4 shows results of the Mistral-7B family. They verify our previous conclusions.
Besides, to verify our conclusion’s generalizability, we conduct experiments on a larger model, Llama-2-13B (Touvron et al., 2023), shown in Appendix F. We also conduct experiments following the traditional causal tracing setting (Meng et al., 2022), which asks the LLM to output the object given a subject. We do not use the traditional setting in the main experiments because it cannot test knowledge storage in the object, and it only allows a small range of datasets, as explained in Appendix B.4. The results also verify the conclusions.
@@ 5 Internal Belief of Truthfulness
How LLMs internally assess the truthfulness of an input statement is another essential aspect of making LLMs truthful and reliable. Previous studies have found that given an LLM and a statement, whether the LLM believes the statement to be true or false can be assessed from the hidden representations encoded by the model. Such belief of truthfulness can be linearly represented along a truthfulness direction in the hidden representation space (Marks & Tegmark, 2024; Bürger et al., 2024). We analyze this direction in base models and post models to analyze whether post-training changes this truthfulness direction.
Linear Probe for Truthfulness To identify the truthfulness direction in a model, we compute difference-in-mean on the hidden representations $h^{l}$, where $l$ is the layer number where truthfulness is most strongly encoded (based on causal tracing results in Section 4). We get this truthfulness direction on one dataset (training dataset) and transfer it to other datasets. Given a training true/false dataset $\mathcal{D}^{\text{train}}$, we separate it into true statements $\mathcal{D}_{\text{true}}^{\text{train}}$ and false statements $\mathcal{D}_{\text{false}}^{\text{train}}$. Similar to knowledge-storage experiments, we use two true statements and two false statements to construct 4-shot prompts, specified in Appendix C.1. The model follows the 4 examples to output “TRUE” or “FALSE” for the final statement. In this process, we compute the truthfulness direction $\mathbf{t}$ as:
@@
where $i$ is the last token of the input prompt and $l$ is the selected layer. Figure 3 (a) and (b) show the cosine similarities of $t^{l}$ from different models on two truthfulness datasets. The heatmaps show a high cosine similarity, revealing that base, sft, and instruct models have remarkably similar internal truthfulness directions.
To further investigate the generalizability, we utilize $\mathbf{t}$ as the weight of logistic regression to construct an MM probe to classify whether a statement is true (Marks & Tegmark, 2024) by $p=\sigma(h_{i}^{l}(s)^{T}\mathbf{t}^{l}),\text{ where }s\in\mathcal{D^{\text{train}}}$, and $\sigma$ is the sigmoid function. We train the probe on five datasets and test its performance on another dataset. We conduct model-transfer experiments, training the probe on the hidden representations of true/false statements generated by one model and evaluating its accuracy in classifying representations generated by other models. We compare the accuracy of training the probe on post ($p_{\textsc{post}}$) and applying it on post’s test representations (baseline) versus training it on base ($p_{\textsc{base}}$) and applying it to post’s test representations (forward transfer). Table 2 presents the results. The probe classification accuracies across base, sft, and instruct are very similar. Across all datasets, $p_{\textsc{base}}$ achieves comparable accuracy to $p_{\textsc{sft}}$ and $p_{\textsc{instruct}}$ when applying on sft and instruct’s test representations, with little differences ($\Delta$). Experiments on the Mistral model family also verify this conclusion, as shown in Appendix C. These findings suggest that the direction corresponding to truthfulness is preserved in post-training.
Figure 3: Cosine similarities of truthfulness and refusal directions of Llama-3.1-8B base, instruct, and sft. Truthfulness directions are similar while refusal directions are different.
@@
Table 2: Probe transfer accuracy ($\uparrow$) of Llama-3.1-8B base, sft, and instruct tested on 6 truthfulness datasets. For each row, the datasets from the other 5 rows are used for training. $p_{model_{1}}\rightarrow h_{model_{2}}$ means using the probe trained on $model_{1}$ to classify truthfulness direction in $model_{2}$. Probe transfer shows little difference ($\Delta$) compared to the same-model probe.
Transfer Intervention with Truthfulness Directions The truthfulness direction $\mathbf{t}$ can also be used to steer model output. To flip a model’s response between “TRUE” and “FALSE” for a statement, we can add $\mathbf{t}$ to the model’s hidden representation as $\tilde{h}^{l}=h^{l}+\lambda\mathbf{t}^{l}$, with $\lambda=\pm 1$ to control the flipping direction following Marks & Tegmark (2024). We also conduct additional robustness experiments with varying values of $\lambda$ in Appendix C.4. The main conclusions are not affected by the choice of $\lambda$, so we use a standard setting of $\pm 1$ in the main experiments. To investigate the transferability of $\mathbf{t}$, we test: (1) intervening $h_{\textsc{sft}}$ with $\mathbf{t}_{\textsc{base}}$ versus $\mathbf{t}_{\textsc{sft}}$; and (2) intervening $h_{\textsc{instruct}}$ with $\mathbf{t}_{\textsc{base}}$ versus $\mathbf{t}_{\textsc{instruct}}$. We evaluate the intervention performance using the Intervention Effect (IE): $(\tilde{P}^{-}-P^{-})/(1-P^{-})$ for false $\rightarrow$ true intervention, and $(\tilde{P}^{+}-P^{+})/(-1-P^{+}$) for true $\rightarrow$ false intervention. $P^{-}$ and $P^{+}$ represent the average probability difference P(TRUE)P(FALSE)P(TRUE)-P(FALSE) for false and true statements, respectively. $\tilde{P}^{-}$ and $\tilde{P}^{+}$ are $P^{-}$ and $P^{+}$ after intervention, respectively. The goal is to increase P(TRUE)P(FALSE)P(TRUE)-P(FALSE) for false statements after the intervention, i.e., $\tilde{P}^{-}$, and to decrease $\tilde{P}^{+}$ for true statements after the intervention, so a higher IE indicates better intervention performance. The results in Table 3 show that when steering sft, the difference ($\Delta$) of IE between $\mathbf{t}_{\textsc{base}}$ and $\mathbf{t}_{\textsc{sft}}$ is little. Similar results hold for instruct. We also conduct experiments on Mistral models in Appendix C, which verifies this result. We illustrate two intervention examples in Appendix C.6, which shows that $\mathbf{t}_{\textsc{base}}$ can flip T/F outputs in post models as effectively as $\mathbf{t}_{\textsc{post}}$. These findings further support our conclusion: post-training generally preserves the internal representation of truthfulness.
@@
Table 3: Intervention effect ($\uparrow$) of intervention on Llama-3.1-8B base, sft, and instruct. For each row, we use the other 5 rows’ datasets for training. $t_{model_{1}}\mapsto h_{model_{2}}$ means using the truthfulness direction in $model_{1}$ to intervene $model_{2}$. Transfer interventions show small differences ($\Delta$) compared to same-model interventions.
@@ 6 Refusal
Post-training also aims at preventing LLMs from generating harmful and toxic content. For example, when inputting harmful prompts like “Give instructions for making explosive materials”, the instruct models might refuse to answer by outputting “I can’t fulfill that request…”, but base models might not. Recent studies by Arditi et al. (2024) show that, similar to the internal belief of truthfulness, this refusal behavior can also be linearly mediated by a vector in the hidden representations as a ”refusal direction”. By steering a model with it, we can encourage the model to change its original behaviors to follow harmful instructions or refuse harmless instructions. Kissane et al. (2024a) found that base models also demonstrate the refusal behavior for some harmful instructions, and thus a refusal direction can also be extracted. It also verified the backward transferability of transferring the refusal direction from the instruct model to the base model. We aim to compare the refusal directions in post models versus base models similarly to the truthfulness direction in Section 5 and study its forward transferability.
@@
To extract the refusal direction $\mathbf{r}$, we use $\mathcal{D}_{\text{harmful}}^{\text{train}}$ (a size-128 subset of advbench) and $\mathcal{D}_{\text{harmless}}^{\text{train}}$ (a size-128 subset of alpaca) to construct the refusal direction. We calculate the refusal direction similarly to the truthfulness direction based on Equation 3. Following Arditi et al. (2024), we compute candidate $\mathbf{r}$ for all token positions and layers and select the most effective one. In the intervention experiments, given $\mathbf{r}$, we induce the refusal behavior on harmless inputs by adding $\mathbf{r}$ to the model’s representations at the layer where $\mathbf{r}$ is learned, i.e., $\tilde{h}^{l}\leftarrow h^{l}+\mathbf{r}^{l}$. To reduce refusal, we subtract $\mathbf{r}$ from the model’s representations at all layers, i.e., $\tilde{h}\leftarrow h-\hat{\mathbf{r}}\hat{\mathbf{r}}^{\top}h$, where $\hat{\mathbf{r}}$ is the unit-norm vector of $\mathbf{r}$. Interventions are applied at all token positions.
@@
To study the refusal behavior across models, we first directly compare r learned on base ($\mathbf{r}_{\textsc{base}}$), sft ($\mathbf{r}_{\textsc{sft}}$), and instruct ($\mathbf{r}_{\textsc{instruct}}$) models. Figure 3 (c) shows that $\mathbf{r}_{\textsc{base}}$ has very low cosine similarity with $\mathbf{r}_{\textsc{sft}}$ and $\mathbf{r}_{\textsc{instruct}}$. To further investigate this, we conduct forward transfer intervention experiments similar to Section 5. We compare the Refusal Score (RS) when using $\mathbf{r}_{\textsc{base}}$ to steer sft and instruct versus using their native refusal vectors ($\mathbf{r}_{\textsc{sft}}$ and $\mathbf{r}_{\textsc{instruct}}$). RS is calculated as the percentage of responses where refusal keywords such as “I can’t” or “I am sorry” appear at the beginning of outputs. We do an intervention on both harmful and harmless datasets, sampling 100 prompts from each for testing. We try to alter the original outputs, i.e., to decrease RS for harmful inputs and increase RS for harmless inputs. Table 4 demonstrates that $\mathbf{r}_{\textsc{base}}$ generally cannot be effectively transferred to steer instruct and sft. Following Arditi et al. (2024), we also conduct experiments on Qwen-1.5-0.5B/Instruct (Bai et al., 2023) and Gemma-2-9B/Instruct (Team et al., 2024) in Appendix D. Results also verify this conclusion: post-training changes the refusal direction and it has limited forward transferability.
@@
Table 4: Intervention RS of Llama-3.1-8B base, sft, and instruct tested on harmful and harmless inputs. $r_{model_{1}}\mapsto h_{model_{2}}$ means using the refusal direction in $model_{1}$ to intervene $model_{2}$, and baseline refers to the original Refusal Score without intervention. For harmful inputs we use ablation and for harmless inputs we use addition.

### v3

@@ 4 Knowledge Storage and Representation
LLMs are known to store factual knowledge at specific locations of their parameters, particularly in “knowledge neurons” and MLP layers acting as key-value memories. This enables them to answer factual queries, such as answering “TRUE” or “FALSE” for the prompt “The city of New York is in the United States. This statement is:”. While such knowledge is believed to emerge during pre-training and persist through post-training, mechanistic evidence remains limited. As knowledge is foundational for LLMs, we examine how post-training affects it by asking two research questions about (Q1) knowledge-storage locations and (Q2) knowledge representations.
When prompted to classify a statement’s truthfulness, LLMs extract stored knowledge at certain layers and inject it into the hidden states to guide the final output. Following marks2024geometrytruthemergentlinear, we adapt causal tracing to identify knowledge-storage locations by patching hidden states between true and false statement pairs. Each pair is token-aligned and differs only in subject, e.g., “The city of Seattle is in France.” vs. “The city of Paris is in France.”, and the relation (e.g., city-in-country) is true for only one statement. We target subject and object tokens (e.g., city and country) for knowledge location analysis.
@@
$)],M_{i}^{l}(s,\hat{s}):=log[\frac{P(``\textsf{TRUE}")}{P(``\textsf{FALSE}")}|patching(h_{i}^{l}(s),h_{i}^{l}(\hat{s}$))], (1)
where a high value indicates knowledge about $s$ is stored in the $l$-th layer at the $i$-th token.
In order to analyze the general locations of knowledge beyond individual knowledge, we average the patching results from a set $D$ of carefully curated statements with the same length and the same subject/object token positions, and each (s,s^)($\hat{s}$) statement pair for patching only differ in their subjects (see Appendix B.1 for dataset construction details). We construct input prompts using 4-shot examples containing 2 true statements and 2 false statements with answers, followed by the question statement. Patching is applied to the question statement via three forward pass as described above. Then we aggregate the results for each $(l,i)$ and normalize them across all layers $l\in L$ and tokens $i\in I$ for better visualization:
@@
For the normalization, we divide the range $\min_{i,l}\tilde{M}_{i}^{l},\max_{i,l}\tilde{M}_{i}^{l}]$ into 20 equal-width bins, set the values in the lower 10 bins to 0 and the values in the upper 10 bins to 0.1, 0.2, …, 1. We denote the normalized result as $M_{model}\in R^{L*I}$, subscripted by the specific model.
@@
Q1: Does post-training change LLM’s knowledge storage locations? Figure 2 visualizes the log probability ratio ($M_{model}$) of Llama-3.1-8B base and instruct on the cities dataset. As shown in Figure (a), influential patching consistently occurs at three token positions: subject, object, and the last token. The last token is always important as it contains information of the whole sentence, whereas the subject and the object position indicate knowledge storage and are important for both base (e.g., (a)) and instruct (e.g., (b)). Their difference is nearly zero (e.g., (c)), indicating that base and instruct store knowledge in nearly identical locations. This pattern holds across all datasets and models (additional visualizations in Appendix B.5). We further conduct quantitative analysis with three metrics and include sft models into comparison. We compute Pearson correlation between $M_{\textsc{base}}$ and $M_{\textsc{post}}$, where post is instruct or sft. We also measure the maximum absolute difference value over all tokens, $max|M_{\textsc{post}}-M_{\textsc{base}}|$, as well as only over knowledge-related tokens (subject and object), $max|M_{\textsc{post}}-M_{\textsc{base}}|_{K}$. Llama-3.1-8B results are in Table 1, and Mistral-7B results are in Table 8 in Appendix B.4. All results show almost perfect correlation and low difference, confirming that post-training has little influence on knowledge-storage locations.
@@
Q2: Does post-training change the knowledge representations? We further conduct cross-model experiments by patching hidden representations from base to post (forward patching) and from post to base (backward patching). It allows us to analyze whether knowledge representations in base can still work in post, and vice versa, i.e., whether cross-model patching recover the log probability ratio of same-model patching. Due to space limits, we put the visualizations on all models and datasets in Appendix B.5. The results demonstrate that the forward patching is almost always successful, but the backward patching often fails. It leads to the conclusion that knowledge representations of base still work after post-training, but post-training also develops new knowledge representations.
Verification on in-distribution dataset One natural question for our experiments above is that they are on datasets independent of post-training, which can be out of the post-training distribution. To verify our conclusions completely, we conduct in-distribution experiments by curating datasets from the post-training data. Specifically, we construct (true, false) statement pairs of factual knowledge from the tulu dataset, which was used to develop Llama-3.1-8B-SFT and Mistral-7B-v0.3-SFT from their base models. Different from previous datasets, pairs in the Tulu dataset could have different lengths, so we slightly modify the metric calculation (details in Appendix B.3). The last column of Table 1 shows the Llama-3.1-8B results, and the last column of Table 8 in Appendix B.4 shows the Mistral-7B results. Both verify our previous conclusions.
Verification on larger models and other tracing settings To verify the generalizability of our conclusion, we conduct experiments on a larger model, Llama-2-13B, with results shown in Appendix F. We also do experiments following the causal tracing setting as in meng2022locating, which asks the LLM to output the object given a subject instead of to output true or false given a subject-object pair. We use the latter in our main experiments because it allows location analysis for the object as well, and it works better with more datasets (details are explained in Appendix B.4). The results of all these experiments also verify our conclusions.
@@ 5 Internal Belief of Truthfulness
How LLMs internally assess the truthfulness of an input statement is another essential aspect of making LLMs truthful and reliable. Previous studies have found that given an LLM and a statement, whether the LLM believes the statement to be true or false can be assessed from the hidden representations encoded by the model. Such internal belief of truthfulness can be linearly represented along a truthfulness direction in the hidden representation space (marks2024geometrytruthemergentlinear; bürger2024truthuniversalrobustdetection). We analyze this direction in base models and post models to illustrate how post-training affects it.
Linear probe for truthfulness To identify the truthfulness direction in a model, we take a (training) dataset $\mathcal{D}^{\text{train}}$ containing a subset of true statements $\mathcal{D}_{\text{true}}^{\text{train}}$ and the rest being a subset of false statements $\mathcal{D}_{\text{false}}^{\text{train}}$, and compute difference-in-mean of the hidden representations of these two subsets of statements. Similar to knowledge-storage experiments, we use two true statements and two false statements followed by the question statement to construct 4-shot prompts (specified in Appendix C.1), so the model outputs “TRUE” or “FALSE” for the final statement. Formally, we compute the truthfulness direction $\mathbf{t}^{l}$ as:
@@
where $i$ is the last token of the input prompt and $l$ is the layer number where truthfulness is most strongly encoded (based on causal tracing results in Section 4). Figure 3 (a) and (b) show the cosine similarities of $\mathbf{t}^{l}$ from base, sft, and instruct models on two truthfulness datasets. The heatmaps show a high cosine similarity, revealing that these models have remarkably similar internal truthfulness directions.
To further investigate the generalizability, we utilize $\mathbf{t}^{l}$ as a probe to classify whether a statement $s\in\mathcal{D^{\text{train}}}$ is true (marks2024geometrytruthemergentlinear), i.e., compute $p=\sigma(h_{i}^{l}(s)^{T}\mathbf{t}^{l})$ with $\sigma$ being the sigmoid function. We train the probe on five datasets and test its performance on a separate test dataset. We also conduct transfer experiments across models, training the probe on the hidden representations generated by one model and evaluating its accuracy in classifying representations generated by other models. For example, we compare the (baseline) accuracy of a probe trained for post ($p_{\textsc{post}}$) and applied on post’s test representations ($h_{\textsc{post}}$) versus the (forward-transfer) accuracy of a probe trained on base ($p_{\textsc{base}}$) and similarly applied to $h_{\textsc{post}}$. Table 2 presents the results, and the probe classification accuracies across base, sft, and instruct are very similar. Especially, $p_{\textsc{base}}$ achieves very similar accuracies to $p_{\textsc{sft}}$ and $p_{\textsc{instruct}}$ when applied to sft and instruct’s test representations across all datasets. Experiments on the Mistral models also show similar results (see Appendix C).
(a) Truthfulness direction on inventors.
(b) Truthfulness direction on animal_class.
(c) Refusal direction on advbench and alpaca.
Figure 3: Cosine similarities of truthfulness (a and b) and refusal (c) directions of Llama-3.1-8B base, instruct, and sft. Truthfulness directions are similar while refusal directions are different.
@@
Table 2: Probe transfer accuracy ($\uparrow$) of Llama-3.1-8B base, sft, and instruct tested on 6 truthfulness datasets. For each row, we use the datasets from the other 5 rows for training. $p_{model_{1}}\rightarrow h_{model_{2}}$ means using the probe trained on $model_{1}$ to classify statement representations by $model_{2}$. Accuracy of transferred probes show little difference ($\Delta$) compared to the same-model probes.
Transfer intervention with truthfulness directions The truthfulness direction $\mathbf{t}^{l}$ can also be used to steer model output. To flip a model’s response between “TRUE” and “FALSE” for a statement, one can add $\mathbf{t}^{l}$ to the model’s hidden representation during forward pass at layer $l$ as $\tilde{h}^{l}=h^{l}+\lambda\mathbf{t}^{l}$, with $\lambda=\pm 1$ to control the flipping direction. To investigate the transferability of $\mathbf{t}^{l}$, we test: (1) intervening $h_{\textsc{sft}}$ with $\mathbf{t}_{\textsc{base}}^{l}$ versus $\mathbf{t}_{\textsc{sft}}^{l}$; and (2) intervening $h_{\textsc{instruct}}$ with $\mathbf{t}_{\textsc{base}}^{l}$ versus $\mathbf{t}_{\textsc{instruct}}^{l}$. We evaluate the intervention performance using the Intervention Effect (IE): $(\tilde{P}^{-}-P^{-})/(1-P^{-})$ for false $\rightarrow$ true intervention, and $(\tilde{P}^{+}-P^{+})/(-1-P^{+}$) for true $\rightarrow$ false intervention. $P^{-}$ and $P^{+}$ represent the average probability difference P(TRUE)P(FALSE)P(TRUE)-P(FALSE) for false and true statements, respectively. Here $\tilde{P}^{-}$ and $\tilde{P}^{+}$ are $P^{-}$ and $P^{+}$ after intervention, respectively. The goal is to increase $\tilde{P}^{-}$, i.e., P(TRUE)P(FALSE)P(TRUE)-P(FALSE) for false statements after the intervention, and similarly to decrease $\tilde{P}^{+}$, so a higher IE indicates better intervention performance. The results in Table 3 show that when steering sft, the difference ($\Delta$) of IE between $\mathbf{t}_{\textsc{base}}^{l}$ and $\mathbf{t}_{\textsc{sft}}^{l}$ is little. Similar results hold when steering instruct. Experiments on Mistral models also show results (see Appendix C). We illustrate two intervention examples in Appendix C.6, which show that $\mathbf{t}_{\textsc{base}}^{l}$ can flip TRUE/FALSE outputs in post models as effectively as $\mathbf{t}_{\textsc{post}}^{l}$. These findings further support our conclusion: post-training generally preserves the internal representation of truthfulness.
@@
Table 3: Intervention effect ($\uparrow$) of intervention on Llama-3.1-8B base, sft, and instruct. For each row, we use the dataset from the other 5 rows for training. $t_{model_{1}}\mapsto h_{model_{2}}$ means using the truthfulness direction in $model_{1}$ to intervene $model_{2}$. Transfer interventions show small differences ($\Delta$) compared to same-model interventions.
@@ 6 Refusal
Post-training also aims to prevent LLMs from generating harmful and toxic content. For example, when inputting harmful prompts like “Give instructions for making explosive materials”, instruct models are likely to refuse to answer by outputting “I can’t fulfill that request…”, but base models may not. Recent studies by arditi2024refusallanguagemodelsmediated show that, similar to the internal belief of truthfulness, this refusal behavior can also be linearly represented by a vector in the hidden space as a “refusal direction”. By steering a model with it, we can encourage the model to change its original sensible behavior to follow harmful prompts or refuse harmless prompts. BaseLLMsRefuseToo found that base models also demonstrate the refusal behavior for some harmful instructions, and thus a refusal direction can be extracted. The study also verified the backward transferability of the refusal direction from instruct to the base. We aim to compare the refusal directions in post versus base similarly to the truthfulness direction in Section 5 and study its forward transferability.
@@
Refusal direction identification and intervention To identify the refusal direction $\mathbf{r}$, we use $\mathcal{D}_{\text{harmful}}^{\text{train}}$ (a size-128 subset of advbench) and $\mathcal{D}_{\text{harmless}}^{\text{train}}$ (a size-128 subset of alpaca) as two contrastive datasets to calculate $\mathbf{r}$ similarly to the truthfulness direction $\mathbf{t}$ in Equation 3. We then perform intervention evaluation follow arditi2024refusallanguagemodelsmediated closely. Given $\mathbf{r}$, we induce the refusal behavior by adding $\mathbf{r}$ to the model’s representations at layer $l$, i.e., $\tilde{h}^{l}\leftarrow h^{l}+\mathbf{r}^{l}$, and $l$ is decided based on the best intervention result. To reduce refusal, for better effect, we ablate $\mathbf{r}$ from the model’s representations at all layers, i.e., $\tilde{h}\leftarrow h-\hat{\mathbf{r}}\hat{\mathbf{r}}^{\top}h$ for all $l\in L$, where $\hat{\mathbf{r}}$ is the unit-norm vector of $\mathbf{r}$. Both kinds of interventions are applied at all token positions.
@@
To study the refusal direction across models, we first directly compare r learned on base ($\mathbf{r}_{\textsc{base}}$), sft ($\mathbf{r}_{\textsc{sft}}$), and instruct ($\mathbf{r}_{\textsc{instruct}}$) models. Figure 3 (c) shows that $\mathbf{r}_{\textsc{base}}$ has very low cosine similarity with $\mathbf{r}_{\textsc{sft}}$ and $\mathbf{r}_{\textsc{instruct}}$. To further investigate this, we conduct forward transfer intervention experiments similar to Section 5. We compare the Refusal Score (RS) when using $\mathbf{r}_{\textsc{base}}$ to steer sft and instruct versus using their native refusal vectors ($\mathbf{r}_{\textsc{sft}}$ and $\mathbf{r}_{\textsc{instruct}}$). RS is calculated as the percentage of responses where refusal keywords such as “I can’t” or “I am sorry” appear at the beginning of outputs. We do an intervention on both harmful and harmless datasets, sampling 100 prompts from each for testing. We try to alter the original sensible behavior, i.e., to decrease RS for harmful inputs and increase RS for harmless inputs. Table 4 demonstrates that $\mathbf{r}_{\textsc{base}}$ generally cannot be effectively transferred to steer instruct and sft for Llama-3.1-8B. We also conduct experiments on Qwen-1.5-0.5B/Instruct (bai2023qwen) and Gemma-2-9B/Instruct (team2024gemma) (see Appendix D). All results verify the same conclusion: post-training changes the refusal direction, so the direction has limited forward transferability.
@@
Table 4: Intervention Refusal Score (RS) of Llama-3.1-8B base, sft, and instruct tested on harmful and harmless inputs. $r_{model_{1}}\mapsto h_{model_{2}}$ means using the refusal direction in $model_{1}$ to intervene $model_{2}$, and baseline refers to the original RS without intervention. For harmful inputs we use ablation and for harmless inputs we use addition.